Two metal spheres A and B of radius r and 2r whose
centres are separated by a distance of 6r are given
charge Q, are at potential V1 and V2. Find the ratio of
V1/V2. These spheres are connected to each other with
the help of a connecting wire keeping the separation
unchanged, what is the amount of charge that will flow
through the wire?

Grade:Upto college level

3 Answers

Vikas TU
6 years ago
Dear Student,
VA=V1=kq/r
VB=V2=kq/2r
V1/V2= 2/1
This would be the ratio.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Elvis Sam
3 years ago
Dear Student,
 
V= V= kq/r + kq/6r
             = 7kq/6r
 
V= V= kq/2r + kq/6r
             = 2kq/3r
 
V1/V2 = 7/4
 
Ratio = 7:4
 
 
saransh jain
3 years ago
Here,
 
VA=V1=kq/r

VB=V2=kq/2r

V1/V2= 2/1
 
so the ratio = 2:1
 
This would be the ratio.

Regards 
Saransh Jain